OpenStudy (anonymous):

The radioactive isotope thallium-201 with a half life of 73 hours determine the decay rate @hartnn and @ranga

OpenStudy (ranga):

Decay constant / rate = ln(2) / half life

OpenStudy (anonymous):

So whatt would my constant be ?

OpenStudy (anonymous):

So ln(2)/73 ??

OpenStudy (ranga):

yes. Decay constant = ln(2)/73 per hour or you can just leave it as a number with no units attached.

OpenStudy (ranga):

\[\Large N = N _{0}e^{-\lambda t}\]N is the amount of the isotope at time t. \[ N _{0}\]is the amount of isotope at the start and \[\lambda\]is the decay constant. Half life is the value of t when \[N / N _{0} = 1/2\]\[\Large 1/2 = e^{-\lambda * 73}\]Solve for \[\lambda\]\[\lambda = ln(2) / 73\]